# Benchmarking Resilience and Sensitivity of Polyurethane-Based Vision-Based Tactile Sensors

### v1

@@ VII CONCLUSION
The practical deployment of vision-based tactile sensors in real-world environments is often hindered by a lack of physical sensor resilience, as their silicone gels are prone to tearing, abrasion, and delamination. In this work, we introduce a polyurethane-based gel as a resilient alternative and propose a set of standardized, model-free evaluations to compare its resilience and sensitivity to its silicone counterparts.
The results of our resilience evaluations show that polyurethane gels can indeed outlast silicone gels across repeated cycles of compression, shear, and abrasion, resisting the catastrophic punctures and tears experienced by many of the silicone samples. While our model-free sensitivity analysis confirms that silicone provides enhanced force and spatial sensitivity compared to polyurethane at low load, it also reveals that polyurethane can provide high spatial sensitivity and better force sensitivity with increased load.
The use of silicone or polyurethane gels should depend on the application. For delicate manipulation tasks where high sensitivity to forces and contact geometries are the priority, silicone remains a suitable option. However, for applications where sensors must be deployed into different unstructured environments without easy access to replacements, we suggest that polyurethane gels offer an advantage. But, we test on one specific version of each material, and many silicones and polyurethanes exist with different mechanical properties. Further exploration of materials and fabrication techniques may yield better solutions, and we propose that our resilience and sensitivity evaluation techniques provide simple, consistent ways to compare across them.
@@ ACKNOWLEDGMENT
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. The authors would like to thank Roberto Calandra for his guidance throughout the project, and Gia Jeong for her help with the demonstration. The authors also acknowledge the support of the members of the Embodied Dexterity Group.

### v2

@@ VII CONCLUSION
The practical deployment of vision-based tactile sensors in real-world environments is often hindered by a lack of physical sensor resilience, as their silicone gels are prone to tearing, abrasion, and delamination. In this work, we introduce polyurethane gels as more resilient alternatives and propose a set of repeatable, learning-free evaluations to compare their resilience and sensitivity to a representative silicone baseline.
The results of our resilience evaluations show that polyurethane gels can indeed outlast silicone gels across repeated cycles of compression, shear, and abrasion, delaying or resisting the catastrophic punctures and tears experienced by many of the silicone samples. While our learning-free sensitivity analysis confirms that silicone provides enhanced force and spatial sensitivity compared to polyurethane at low load, it also reveals that polyurethane can provide high spatial sensitivity and better force sensitivity with increased load.
The use of silicone or polyurethane gels should depend on the application. For delicate manipulation tasks where high sensitivity to forces and contact geometries are the priority, silicone remains a suitable option. However, for applications where sensors must be deployed into different unstructured environments without easy access to replacements, we suggest that polyurethane gels offer an advantage that scales with hardness. However, our study is limited to three gel material configurations. Further exploration of materials and fabrication techniques may yield better solutions, and we propose that our resilience and sensitivity evaluation techniques provide simple, consistent ways to compare across them.
@@ ACKNOWLEDGMENT
This material is based upon work supported by Google and the National Science Foundation Graduate Research Fellowship Program under Grant No. 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. The authors would like to thank Roberto Calandra for his guidance throughout the project, and Gia Jeong for her help with the demonstration. The authors also acknowledge the support of the members of the Embodied Dexterity Group.